### v2

# Large Language Models, scientific knowledge and factuality: A systematic analysis in antibiotic discovery

Abstract
Inferring over and extracting information from Large Language Models (LLMs) trained on a large corpus of scientific literature can potentially drive a new era in biomedical research, reducing the barriers for accessing existing medical evidence. This work examines the potential of LLMs for dialoguing with biomedical background knowledge, using the context of antibiotic discovery. The systematic analysis is applied to ten state-of-the-art models, from models specialised on biomedical scientific corpora to general models such as ChatGPT, GPT-4 and Llama 2 in two prompting-based tasks: chemical compound definition generation and chemical compound-fungus relation determination. The work provides a systematic assessment on the ability of LLMs to encode and express these relations, verifying for fluency, prompt-alignment, semantic coherence, factual knowledge and specificity of generated responses. Results show that while recent models have improved in fluency, factual accuracy is still low and models are biased towards over-represented entities. The ability of LLMs to serve as biomedical knowledge bases is questioned, and the need for additional systematic evaluation frameworks is highlighted. The best performing GPT-4 produced a factual definition for 70% of chemical compounds and 43.6% factual relations to fungi, whereas the best open source model BioGPT-large 30% of the compounds and 30% of the relations for the best-performing prompt. The results show that while LLMs are currently not fit for purpose to be used as biomedical factual knowledge bases, there is a promising emerging property in the direction of factuality as the models become domain specialised, scale-up in size and level of human feedback.
1 Introduction
Recently, artificial intelligence (AI) algorithms have significantly facilitated the acceleration of antibiotic discovery through high-throughput drug screening, with a particular emphasis on the identification of novel antibiotic compounds [1]. Deep learning (DL) methods have proven effective in isolating small molecule antibiotics exhibiting potent activity against a diverse array of bacterial pathogens [2]. This advancement in machine learning (ML) for drug discovery has been bolstered by the widespread accessibility of datasets emerging from empirical analyses. Notably, the discovery of novel compounds primarily relies on protein structure prediction and modeling of antibiotic targets. Moreover, since the knowledge related to antibiotic discovery is communicated as text in the scientific literature, patents and clinical trials, this creates an opportunity for Large Language Models (LLMs), especially Generative Pre-trained Transformers (GPTs), to serve as proxy for background scientific knowledge, reducing the barriers for integrating and mapping the landscape of available evidence. Eliciting information from LLMs trained on large corpora of scientific, biomedical literature could potentially support the antibiotic discovery process.
Large Language Models have evolved in the past two years significantly influencing natural language processing (NLP) with a variable level of adoption in different specific domains [3, 4]. GPT-1 (Generative Pre-trained Transformer) [5], the first LLM trained using self-supervised pre-training, predicts the next word in a sequence given the previous words, performing this task unidirectionally from left to right. This enables GPT-1 to learn the dependency relations between words and generate relevant and contextual content. GPT-2 [6] was trained with a larger architecture on a larger and more diverse corpus than GPT-1, with 1.5 billion parameters, thus demonstrating better generalisation properties.
GPT-3 [7] with 175 billion parameters represents a phase transition in the evolutionary arch of language models, where the in-context learning properties were noticeably present. Subsequently, InstructGPT [8] extended this foundation, using reinforcement learning from human feedback (RLHF) as a training strategy, leveraging on the ability of the pre-training model to re-align at instruction/intensional level. ChatGPT [9] follows the same architectural principles of InstructGPT, using GPT-3.5 (both larger corpora, and human-feedback prompting datasets). Despite having significantly less domain-specific biomedical text, it can still deliver the ability to encode and articulate more specialised discourse, for example, achieving a score equivalent to a third-year medical student [10].
Despite their comprehensive linguistic interpretation properties, LLMs have inherent operational properties, most notably hallucinations [11, 12], i.e. divergence from factual knowledge and spurious inference, despite its fluency and syntactic correctness [13, 14]. The recently released GPT-4 is not free from this limitation [15]. Hallucinations tend to surface prominently in domain-specific discourse, particularly in biomedical sciences. This elicits the need to establish the ability of these models to encode and preserve factual knowledge as well as establishing methodologies for the critical evaluation of its representation properties [16, 17, 18].
In a range of benchmarks assessing helpfulness and safety, the one of the latest open-source LLMs, Llama 2, typically outperforms current open-source alternatives. Llama 2 was pretrained on publicly available online data sources (on 2 trillion tokens) and its fine-tuned models have been trained on over 1 million human annotations. As the Llama2’s authors [19] emphasized, testing conducted to date has been in English and has not — and could not — cover all scenarios. None of the scenarios tested so far included the use of the model for the purpose discussed in this study. We evaluated variants of this model with 7B, 13B, and 70B parameters.
Figure 1: The overview of the study. Investigation of the ability of LLMs to encode and access biological factual/relational knowledge. More specifically, a highly relevant domain of application was selected, which is motivated by antibiotic discovery.
In this work we systematically assess the biological relational knowledge encoded in state-of-the-art LLMs, namely: GPT-2, GPT-3, GPT-neo, Bloom, BioGPT, BioGPT-Large, Galactica, ChatGPT, GPT-4 and variants of Llama 2 with 7B, 13B, and 70B parameters. We focus on systematically determining the ability of LLMs to capture fundamental entities (fungi, chemicals and antibiotic properties), their relations and supporting facts, which are fundamental for delivering inference in the context of downstream biomedical inference (Figure 1). For instance, we aim to answer the question on whether LLMs can faithfully capture biological domain-specific factual knowledge such as:
  • “Aspergillus fumigatus produces festuclavine"
  • “Cycloclavine is biosynthesised by Aspergillus japonicus"
  • “Enfumafungin demonstrates antibiotic activity against several bacterial, yeast and fungal species"
  • “Enfumafungin and ergokonin A are isolated from Hormonema and Trichoderma species and demonstrate antibiotic activity against several bacterial, yeast and fungal species"
This overarching question is translated into the following specific research questions (RQs):
  1. RQ1
    Do large language models encode biomedical domain knowledge at entity level (e.g. fungi, chemical, antibiotic properties) and at a relational level?
  2. RQ2
    Are there significant differences between different models in their ability of encoding domain knowledge?
  3. RQ3
    Which model performs better in faithfully encoding biological relations?
  4. RQ4
    How do we evaluate the limitations of LLMs in factual knowledge extraction?
This work targets the following contributions:
  • Introduction of a comprehensive framework tailored to systematically assess LLMs capacity to encapsulate domain-specific factual knowledge, focusing on target entities and their interrelations within the biomedical sphere. This framework is a direct response to the labor-intensive nature of text generation evaluation and aims at reducing this burden
  • We provide a qualitative evaluation and comparison of the state-of-the-art LLMs in the context of biomedical inference and factual knowledge extraction. We broaden the scope of existing research from other fields, catering to the often-expressed demand for more comprehensive evaluations of LLMs across diverse disciplines.
  • We explore the extent to which these models rely on memorization versus genuine comprehension of the intricate symbolic relationships between biomedical terms
  • Finally, we present a compelling case and substantial proof supporting the necessity for evaluation frameworks. These frameworks are crucial for identifying the limitations of LLMs and ensuring the effective evaluation of the text they generate.
This study is organized as follows: first, we define our contribution and link it with the existing frameworks and other investigations within the biomedical domain. Then, in the Methodology section, we present a framework for the qualitative evaluation of the biomedical knowledge embedded in LLMs, followed by a detailed description of the methods. In the Results section we apply the proposed framework to ten LLMs and evaluate them for ten chemical compounds with a detailed description of the generated responses. We then discuss our findings in the broader context of reported LLMs’ limitations. We review related work in the field of LLMs and their knowledge representation properties. Finally, we conclude the paper with a summary of the contributions and with future perspectives on the role of LLMs in biomedical inference.
The workflow of the analysis is depicted in Figure 2, and the proposed evaluation framework in Figure 3.
2 Motivational scenario
Antimicrobial resistance in clinically relevant bacteria compromises the effectiveness of existing antibiotics and call for the discovery of new antibiotic substances. Antibiotic development is a slow, costly and failure-prone process that can take over a decade and costs can reach the scale of hundreds of millions of dollars [20]. In a study of nearly 186,000 clinical trials spanning over 21,000 compounds, the probability of success of new drugs treating infectious diseases was found to be 25.2% [21]. However since 2017, only eight new antibiotics have been approved, including one for the treatment of multidrug-resistant tuberculosis (2017-2021) [22]. Furthermore, antibiotics in clinical trials are generally analogues of existing drugs for which antimicrobial resistance mechanisms have already emerged [23], further highlighting the need for new approaches to antibiotic discovery.
Some organisms, such as fungi or bacteria, naturally produce a variety of secondary metabolites (also called natural products) associated with many biological activities. Penicillin, a beta-lactam antibiotic, is a historical example of fungal metabolites with therapeutic properties. In order to effectively exploit the pool of potential sources of antibiotics, avoiding dead ends and rediscovery, it is essential to identify the secondary metabolites known to be produced by these organisms and which substances have already shown antibiotic activity. This knowledge is mainly communicated in the scientific literature, the size and complexity of which makes it difficult to explore it systematically and at scale.
3 Methodology
In this section we describe the workflow of our analysis (depicted in Figure 2), which consists of the following steps:
  • extraction and curation of a dataset with biological relations
  • definition of text generation tasks for LLMs
  • selection of the target LLMs to be included in the evaluation
  • development of an evaluation framework for the efficient qualitative analysis of generated output, including prompt definition, prompt engineering and evaluation metrics
The above steps allowed us to provide the first comparative analysis of LLMs in the context of tasks related to fungi and antibiotics and is transportable to other biological and biomedical relations. The code and dataset used in the study are available at github [^1] https://github.com/digital-ECMT/LLM_factuality_evaluation_framework.git.
Figure 2: The workflow of the performed analysis. $n_{relations}$ 23 selected chemical compound-fungus pairs; $n_{entity1}$ 10 unique chemical compounds included in those pairs; In prompt engineering: blue - entity 1, orange - added context. Qualitative evaluation performed according the the framework depicted in Fig.3.
3.1 Dataset
The investigation of the biological knowledge stored in LLMs requires a collection of biological relations that are elicited and referred in the scientific literature. Scientific publications are part of the training corpus for most LLMs, thus higher prevalence in literature should lead to better performance in factual knowledge extraction. In our study we created a dataset of biological relations accounting for the amount of supporting evidence (number of associated papers).
Chemical compound-fungus pairs were extracted from the Wikidata Knowledge Graph using the predicate found in taxon (p:P703), along with the number of bibliographic references supporting each relationship (via API). Classification and nomenclature information on fungi were extracted from the Mycobank database [^2] https://www.mycobank.org. Information about antibiotic activity of chemicals were extracted from the ChEBI ontology [^3] https://www.ebi.ac.uk/chebi/downloadsForward.do.
A subset containing only fungal species producing chemicals with antibiotic activity was first created. The pairs with the most bibliographic references were progressively integrated until obtaining 100 distinct subject fungi, which corresponds to 123 pairs. Then, a complementary subset (with an equal number of pairs) involving produced chemicals but without antibiotic activity was added from the same list of fungi. Again, the pairs with the most bibliographic references were selected. Only molecules with a 3-STAR review in ChEBI were considered and only those that have an annotated role in the antimicrobial agent class (CHEBI:33281) were selected for positive examples of antibiotics.
As a result, a dataset with 246 chemical compound-fungus pairs (see Figure 2, N) equally divided between chemicals with and without antibiotic activity was obtained. For the experiments, we randomly selected ten chemicals distributed by antibiotic activity and number of references (five chemicals with and five without antibiotic activity) (see Table S.1). These ten chemicals paired with one or more fungi. In total, there were 23 chemical compound-fungus pairs (see Figure 2, $n_{relations}$) in the analyzed subset. The selected ten chemical compounds were defined as entity 1 for two tasks. The fungus was defined as entity 2 in the second task. Of note, the selected 23 chemical-fungus relations are described by multiple scientific publications in PubMed, all published before 2019 (for exact numbers see Table S.1). This ensures these relations are contained in the LLM training corpus, provided that the corpus contains PubMed articles or abstracts.
3.2 Large Language Models
Table 1: Summary of the evaluated Large Language Models ordered by the release date.
Model Lab Training Corpus Vocabulary size Model size (parameters) used/available Date of the model release* Reference
GPT-2 OpenAI WebText: 40 GB of text, 8M documents, from 45M webpages upvoted on Reddit (all Wikipedia pages removed) 50,257 124M/1.5B Feb 2019 [6]
GPT-3 OpenAI 570 GB plaintext, 0.4 trillion tokens. (410B CommonCrawl, 19B WebText2, 3B English Wikipedia, and two books corpora (12B Books1 and 55B Books2)) 175B June 2020 [7]
GPT-neo EleutherAI The Pile - a large scale curated dataset created by EleutherAI 825GB, 380 billion tokens 50,257 1.3B/2.7B March 2021 [24]
Bloom BigScience 45 natural languages 12 programming languages In 1.5TB of pre-processed text, converted into 350B unique tokens 250,680 560M-3B/176B July 2022 [25]
BioGPT Microsoft 15M PubMed abstracts (trained from scratch) 42,384 347M/347M Oct 2022 [26]
BioGPT-large Microsoft 15M PubMed abstracts (trained from scratch) 42,384 1.5B/1.5B Oct 2022 [26]
Galactica Meta AI 106B tokens; Documents: 48M papers, 2M code, 8M reference material, 2M knowledge bases, 0.9M filtered commoncrawl, 1.3M prompts, 0.02M other 50,000 1.3B/120B Nov 2022 [27]
ChatGPT OpenAI unavailable 20B (175B) Nov 2022
GPT-4 OpenAI unavailable Feb 2023
7B
Llama 2 Meta AI unavailable 13B July 2023 [19]
70B
    *The ‘Date of the model release’ is based on its first arXiv version (if exists) or official public release.
In this study, we evaluated ten LLMs (summarized in Table 1), where six of them were released since July 2022. The models vary in terms of training corpus, vocabulary size, number of parameters, layers and maximum input sequence length. Below, we articulate the essential properties of each target model.
There are two main high-level architectures associated with language models: BERT - Bidirectional Encoder Representations from Transformers (and its derivatives) and GPT - ‘Generative Pre-trained Transformer’ (and its derivatives). Extensive research has been conducted on the former in the biomedical domain, demonstrating considerable success in numerous discriminative downstream tasks [6, 7]. In this paper, we evaluate generative transformer language models pre-trained on large scale general literature (GPT-2, GPT-3, GPT-neo, Galacitca, Bloom, ChatGPT, GPT-4, Llama 2) and specific for the biomedical domain (BioGPT and BioGPT-Large). Due to evaluation in the zero-shot setting, the study focuses on decoder-only architectures due to their superior performance in comparison to encoder-decoder frameworks [28].
GPT is a language generation model that utilizes the transformer decoder architecture for pre-training on large-scale text corpora. GPT is grounded in the next-token prediction task, in which the model learns to predict the next word token based on the preceding word tokens.
GPT-2 adopts GPT-1’s architecture with some differences, including the modified initialization, pre-normalization, and reversible tokenization. A different initialization, accounting for residual path accumulation, scales weights by 1/N, where N is the number of residual layers. Vocabulary expands to 50,257, context size increases to 1024 tokens, and batch size grows to 512. GPT-2 used a larger architecture and a more diverse corpus than GPT-1, including web pages, books, and scientific articles. GPT-3 uses the same model and architecture as GPT-2, with the exception of the use of alternating dense and sparse attention patterns in the layers of the transformer, similar to the Sparse Transformer. Compared to previous generative models, GPT-3 has an order of magnitude larger parameter space (175 billion parameters).
GPT-Neo is an autoregressive transformer language model, which is based on the replication of the GPT-3 architecture. With 2.7 billion parameters, the model was trained on the Pile [24, 29], a large-scale curated corpus containing Pubmed Central and Wikipedia data. The architecture of Bloom (BigScience Large Open-science Open-access Multilingual Language Model [25]) is similar to GPT-3, being trained on 46 different languages and 13 programming languages, including the Pile. The largest Bloom contains 176B parameters, however due to hardware access limitations, we tested Bloom under the 560M and 3B parameter setting.
BioGPT and BioGPT-large are domain-specific generative transformer language models pre-trained on large-scale biomedical literature [30], and are pre-trained on 15M PubMed items (each with both title and abstract) from scratch. We tested both versions: with 347M and 1.5B parameters (BioGPT-large). BioGPT uses the GPT-2 architecture as its backbone. However, instead of relying on the GPT-2 vocabulary, it has created an in-domain vocabulary through Byte-Pair Encoding (BPE) applied to the corpus, leading to a total of 42,384 word pieces. The model utilizes the GPT-2medium as its core network, featuring 24 layers, a hidden size of 1024, and 16 attention heads. This results in a total of 347 million parameters for the BioGPT model.
Galactica is model targeting the parsing of scientific discourse, developing particular tokeneizers for common scientific text (e.g. equations, chemical formulae). It implements the following architectural modifications: GeLU activations [31], 2048-length context window, no biases [32], learned positional embeddings, and 50k-token Byte-Pair Encoding vocabulary [33] from a 2% training data subset. It is available in sizes from 125M to 120B parameters. The largest version that we were able to run on was 1.3B.
Both GPT-4 and ChatGPT builds upon the GPT-3.5 model which is then refined via the Reinforcement Learning with Human Feedback (RLHF) mechanism. Both corpora and human feedback alignment dataset are not reported.
Llama 2 was trained with an extensive corpus of 2 trillion tokens extracted from publicly available sources like Common Crawl (an archive of billions of webpages), Wikipedia, and public domain books from Project Gutenberg. In contrast to the above two models, the training dataset for Llama 2 is ascertainable. Llama 2 exhibits superior performance relative to other open-source language models across diverse external benchmarks, encompassing assessments of logical reasoning, coding proficiency, and general knowledge. However, it is emphasized that this model is just intended to be further trained to tailored to specific requirements. This study evaluates Llama 2 7B, Llama 2 13B and Llama 2 70B.
The final baseline set consists of: GPT-2, GPT-neo (1.3B parameters), BLOOM (3B parameters), BioGPT, BioGPT-large and Galactica via huggingface API [^4] https://huggingface.co; Llama 2 using the llama.cpp [^5] https://github.com/ggerganov/llama.cpp (accessed in October 2023); all above with the default sampling procedure for text generation, i.e. greedy decoding and temperature equal 0; GPT-3 via the OpenAI API; and for ChatGPT and GPT-4 a default OpenAI user interface was used to generate outputs, each prompt in a separate session (accessed in March 2023).
3.3 Framework for the evaluation of Large Language Models on the encoding of factual scientific knowledge
An intrinsic evaluation of the text generated by the models was performed by a human expert according to five criteria: Fluency, Prompt-alignment, Semantic coherence, Factuality and Specificity [34, 35]:
  • Fluency evaluates the quality of the generated text, considering two sub-criteria: syntactic correctness and style. Syntactic correctness assesses whether a sentence conforms to the rules of grammar and expected syntactic structure. Scientific style is determined by the vocabulary of the generated text conforming to a scientific theme (‘reads as a scientific text’).
  • Prompt-alignment refers to the relation of relevance and pragmatic coherence between the prompt and the generated text, whether the generated text is consistent with the input prompt and is related to the entities and context of the prompt.
  • Semantic coherence consists of two levels of analysis: (i) intra-sentence assesses whether the sentence is relating terms surface to expected argument types for a given predicate and (ii) inter-sentence: assessing whether adjacent sentences have a coherent discourse relation [36].
  • Factuality is distinct from the meaning of a sentence which is conveyed by its semantic coherence. A sentence generated by a model is true if there exists a factual statement in the scientific literature, a database, or any other trusted source.
  • Specificity evaluates the alignment between the prompt and the answer regarding their level of abstraction. While some answers can be factually correct, they may not answer to the level of abstraction required by the prompt.
While auto-regressive language models are by design to deliver fluency, prompt-alignment and semantic coherence, text generation during decoding can lead to repetitive, incoherent or meaningless outputs [37]. As they are complementary, the fluency, prompt-alignment and semantic coherence were assessed in one step of the evaluation process (Figure 3, STEP 1). Moreover, the probabilistic nature of the next token prediction task also contrasts with the notion of factuality and large language models are prone to hallucination [38]. Hallucinations are factually incorrect statements which can be difficult to be localised by non-experts due to their expression in a fluent and semantically coherent text. In the context of this study, major hallucinations can be associated with the compound description or relationships with fungi, while minor hallucinations may be associated with the date of first isolation for instance. The factuality of the generated text was assessed in the STEP 2. Importantly, for Relation determination via entity generation, the correctness of the generated entity 2 is verified prior to the verification of the whole description. If incorrect, we assume that the LLM fails to determine the relation. Specificity of the generated text is assessed in the STEP 3. Note that only STEP 1 does not require a domain expert.
Figure 3: The framework for qualitative and efficient evaluation of the Large Language Models in the context of extracting biological relation from these models. The entity 1 stands for chemical compound name. The entity 2 stands for fungus name.
In this study two major tasks were evaluated, described in 2.5. Models were prompted with task-specific prompts (see Table 2) and the generated text was evaluated according to the framework depicted in Figure 3.
3.4 Text generation tasks
We investigate the ability of LLMs to act as knowledge bases and establish the relation between fungi and antibiotics. Following the evaluation framework proposed in section 2.4, we define two text generation tasks (see the workflow in Figure 2).
Task 1. Chemical compound definition generation
Task 1 uses prompts for chemical compound definition generation, shown in the Table 2 as P1-P7. For the same task we define simple prompts (P1, P2) and context-based prompts, elucidating the type of the subject entity: ‘a compound’ or ‘a substance’ (P4 and P5) [39]. For Galactica, according to authors’ suggestions, prompts were framed as questions, rather than sentences to complete (P3, P6, P7). The entity 1 represents the selected chemical compounds. Performance in Task 1 was calculated for each step of the framework (Fig.3) and is shown in Table 3.
Task 2. Chemical-fungus relation determination via entity generation
Task 2 uses prompts for relation determination via entity generation: {entity 1}- {relation} - {entity 2}, where entity 1 is a chemical compound name and entity 2 is a fungus name (P8-P19, see Table 2). The prompts P13-P18 contain the context at the end, guiding the model towards a relation to a fungus, rather than to another organism capable of producing the compound. Evaluation was performed according to the framework (Fig.3) for the entire task, both with and without context. We report the number of responses meeting the criteria of given STEP in the framework for each prompt (x correct out of 110, see Table 4).
Table 2: Prompts used in Task 1 and Task 2
Task Type of prompts No. Prompt
Task 1: Chemical compound definition generation simple prompts P1 {entity 1}
P2 {entity 1} is
[…]
context-based prompts P4 {entity 1} is a compound
P5 {entity 1} is a substance
[…]
Task 2: Chemical-fungus relation determination via entity generation simple prompts P8 {entity 1} is isolated from
P9 {entity 1} originally isolated from
P10 {entity 1} is produced by
P11 {entity 1} is a natural chemical compound found in
P12 {entity 1} is a natural product found in
context-based prompts P13 {entity 1} is isolated from the fungus
P14 {entity 1} is isolated from fungi
P15 {entity 1} is isolated from fungi, such as
P16 {entity 1} is produced by the fungus
P17 {entity 1} is produced by fungi
P18 {entity 1} is produced by fungi, such as
[…]
    *Evaluation for Galactica included P3, P6, P7 and P19 prompts only.
3.5 Expert evaluation
As described in 2.2, we selected ten names of chemical compounds (five chemicals with and five without antibiotic activity). The maximum numbers of tokens to generate, excluding the number of tokens in the prompt is 30, except for ChatGPT, GPT-4 and Llama 2. In total, we investigated outputs for 15 prompts for the models Bloom, GPT-2, GPT-neo, GPT-3, BioGPT, BioGPT-large, ChatGPT, GPT-4 and Llama 2; and four prompts for Galactica. All prompts are presented in Table 2. In both tasks, in order to measure the variance in our small sample size setting, we applied bootstrapping with 1000 iterations, reporting the range (in %) of correct responses for a given prompt and STEP with the 95% confidence interval.
All 1690 outputs generated by the models, i.e. 11 models x 15 prompts x 10 chemical compounds + 1 model (Galactica) x 4 prompts x 10 chemical compounds, were evaluated and labeled according to the proposed framework by 2 reviewers independently of each other, on each of steps 1-3. A total of 1690 records were evaluated on the step 1. 1231 of them were moved to the steps 2 and 3, and then still dually reviewed by two reviewers. During all steps, there were 22 (1.3%) discrepancies, which were resolved by discussion. Two experienced reviewers took part in this systematic analysis. Both reviewers were involved in the study design of the systematic analysis and the definition of the criteria. The reviewers went through a pilot process with the criteria prior to starting the study to ensure they had a similar understanding of the criteria. The evaluation was carried out systematically by two independent experts in the field, one with a PhD in microbiology and chemistry and the other with a PhD in biochemical pharmacology, both having prior experience in the manual assessment and curation of datasets.
4 Results
4.1 Chemical compound definition generation
This task assesses the ability of the LLM to generate a factual definition of a chemical compound (equal entity 1). For instance with the Prompt = ‘{entity 1} is’ with the entity ‘Atromentin’, a correct answer could have the form ‘Atromentin is a natural product…’. We investigated ten examples of chemical compounds (five with and five without antibiotic activity) for the qualitative evaluation according to the previously described criteria (see section 2.4, Table S.1). As prompts, we provided single entities or entities with the more specific context (see section 2.5). The results are summarized in Table 3. Examples of generated answers with the qualitative evaluation are presented in Table S.2. Note, for ChatGPT, GPT-4 and Llama 2 we did not limit the number of tokens and evaluated the complete output. In total, we manually evaluated 470 outputs in this task.
GPT-2 and GPT-neo: no semantic coherence, no factuality. GPT-2 failed to generate any semantically coherent text that is related to the entity and to the target domain (biomedicine in general). After adding ‘is’ to the prompts, the model generates fluent text only for 2/10 compounds, but with no semantic coherence. Adding context leads to even more fluent outputs but still loses semantic coherence and there is no prompt-alignment. For 7/10 compounds very fluent text is generated that is unrelated to either the associated fungus or the antibiotic activity of the compound.
GPT-neo, similarly to GPT-2, failed to generate any text or any fluent text. After adding ‘is’ - the model generates (for 8/10) syntactically correct texts unrelated to the given name, resembling to an n-gram scientific fragment unrelated to the mention. After adding context, the model asserts two compounds (Fumitremorgin C, Conidiogenone) as being produced by fungi (Fusarium sphaerospermum, Penicillium notatum, respectively), however under incorrect associations. The model failed to generate factual knowledge.
GPT-3: Semantically coherent answers less sensitive to prompt design. 3/10 factual answers. Hallucinates fungi names from compounds. GPT-3 generated semantically coherent text regardless of the prompt design for all chemical compounds, both with and without context (see example of Atromentin in Table S.2). It generated a similar number of factual answers regardless of the context (30% vs 35%). Interestingly, adding context to the prompt did not increase specificity (‘Ergosterol is … a sterol found in the cell membranes of fungi. It is a major component of the fungal cell membrane, and is similar to cholesterol in animals. It is an important precursor for the synthesis of other compounds, such as ergocalciferol’ vs ‘Ergosterol is a compound … found in the cell membranes of fungi. It is a sterol, a type of lipid molecule, and is the main sterol in the fungal cell membrane. It is similar to cholesterol, which is found in animal cell membranes’). For 2/10 chemicals (Ergosterol and Alternariol) GPT-3 generated factual and specific answers regardless of the prompt and context. GTP-3 generated incorrect fungi names (in 39.3% of incorrect answers) that consists of the compound name in the prompt, e.g. Myriocin - Streptomyces myriocin, Chloromonicilin - Streptomyces chloromyceticus (Table S.3).
Bloom: no factual knowledge, low semantic coherence: biased towards treatments, repeats the same associated terms or statements. For prompts without context, Bloom fails to generate a syntactically correct output. It does not generate any text (for example, ‘Conidiogenone ’), generates the chemical compound entity as a code within a link (for example, ‘org.apache.thrift.TBase("org.apache.thrift.TBase", "flavasperone",’) or generates text with only special characters (for example, ‘Ergosterol ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñÂâÇ ÊäóÊßÂåñ’). The model tends to repeat the same words or phrases for a given entity (for example, ‘Fumitremorgin C is a potent inhibitor of the enzyme, which is responsible for the synthesis of the enzyme, and is therefore a potent inhibitor of the synthesis of the enzyme.’). In addition, the model generates names that does not exist in the literature (for example, ‘Atromentin is a natural product derived from the fruit of the plant Atrium tridentatum.’). When adding the context to the prompts, Bloom is biased towards considering the compounds as a treatment for a variety of diseases. It generates fluent but no semantically coherent text, unrelated or not specific to the compound. Frequently the answer contains repeated statements and facts, e.g. ‘Fumitremorgin C is a substance that is used in the treatment of cancer. It is a natural product that is used in the treatment of cancer. It is a natural product that is’ - despite its factuality, it corresponds to low prompt-alignment and specificity.
BioGPT: low specificity, adding context leads to even lower specificity, leading to the answer being true for any compound. BioGPT generates fluent and syntactically correct text tends to correspond to a more general discourse type (when contrasted to scientific discourse). For prompts ‘{entity 1}’ and ‘{entity 1} is’, only one compound (Ergosterol) has factual answers. For instance, the generated text states that one compound is an antibiotic, but the rest of the sentence is not related to the compound, or the model confuses DNA polymerase alpha/beta/gamma (Table S.2 for Aphidicolin). When more context is provided (prompt with ‘is a compound’), BioGPT generates a factual text with low specificity for 4/10 compounds. It provides a non-specific, universal definition of a compound, which is referred to regardless of the compound name (Table S.2, example: ‘Fumitremorgin C is a compound with a unique structure and a unique mode of action.’ - the same is for Verrucarin A). After adding the context ‘is a substance’, BioGPT recognises that 5 /10 compounds were produced by fungus, but all incorrect.
BioGPT-large: Semantically very coherent. 3/10 factual answers. Hallucinations under high quality, scientific text are difficult to spot. BioGPT-large generated semantically coherent outputs for 80% of the prompts without and 100% prompts with the context. However, the answers were factual only for 30% and 35% for prompts without and with context, respectively. Despite some of them were not specific, BioGPT-large generated answers better fitting to a scientific discourse level, compared to the previous models (Table S.3). Similarly to GPT-3, for 28% of answers with no factuality, BioGPT-large assigned a wrong producer to the compound, e.g. ’Chloromonilicin is a new antibiotic produced by a strain of Streptomyces. It is active against Gram-positive bacteria and fungi.’ (true answer: is produced by Monilinia fructicola). This is of particular concern, as the generated text imitates correct definitions very well, disguising the lack of factual knowledge. Such hallucination is very difficult to spot and a vigilant expert check is required.
Galactica: poor semantic coherence, improves after adding context to the prompt, answers resembles as corpora fragments. 1/10 factual answers. Galactica only for 2/10 compound names can generate text that is related to the target entity and that is aligned to the domain, but in both cases lacks factuality. For one case, the model assigned that the compound was produced by a fungus but misnamed it (‘Aphidicolin is a polyketide antibiotic produced by the fungus Aspergillus nidulans. It is a potent inhibitor of DNA polymerase alpha.’). Adding the context ‘What is the compound {entity 1}?’ improves the output. The text is more meaningful and fluent, and is domain-aligned. However, it still lacks factual knowledge with one exception for Myriocin, which is recognized as an antibiotic. The generated text resembles the fragment from a title of a specific scientific paper than a general description (‘Myriocin, a new antifungal antibiotic. I. Taxonomy, fermentation, isolation, physico-chemical properties and biological activity.’). Moreover, the model recognises that Fumitremorgin C is a mycotoxin but again misnamed the fungus name (‘Fusarium verticillioides’, instead of ‘Aspergillus fumigatus’).
ChatGPT: Fluent, prompt-alignment and semantically coherent descriptions. Hallucinations are difficult to spot. 6/10 factual answers. ChatGPT generated fluent and semantically coherent text for all prompts. For prompts with no context, the statements were longer and divided into paragraphs (mostly three; note, that a limit was not set on the length of generated answer for ChatGPT). The generated text was coherent regardless of the designed prompts (’{entity 1}’ or ’{entity 1} is’), and 6/10 outputs were factual. The other four generated the wrong fungus name or the wrong chemical formula in the answer. Interestingly, the answers often closely resembled Wikipedia-style descriptions (7/10 compounds - Wikipedia, 2/10 - Wikidata). For prompts with the context, the answers were shorter and mostly consisted of one paragraph. For prompts ’{entity 1} is a compound’ the text contained a chemical formula and a description of the chemical structure. However, in 60% of these prompts the chemical formula was incorrect (see Aphidicolin in Table S.2). The answers that were factual had low specificity, and could satisfy the definition of many other chemical compounds. For prompts ’{entity 1} is a substance’, the answer provides additional biological context and informs about the biological activity of the compound or what it is used against. 70% of these answers contained factual knowledge. Comparing to BioGPT-large, ChatGPT produces even more fluent, prompt-alignment and semantically coherent answers. Due to semantic fluency, evaluation of the factual knowledge is very challenging, and requires both the expert knowledge and time-consuming fact verification. In other words, generated answers can be easily regarded as true and factual for a non-expert, creating a real risk of deriving false facts from the model.
GPT-4: Top of the list and epistemically-aware (admits to lack of knowledge). 14/20 factual answers. For prompts without context, the model gave 70% (14/20) of the answers that were specific and with factuality. In two prompts (out of 20), the characterisation of a given chemical compound was non-factual due to the incorrect name of the fungus that may produce the chemical. Importantly, the model did not guess the definition if it did not recognize the name. For two chemical compounds, the model generated a request for more context or referred to the knowledge cutoff date of 2021-09. After adding context, GPT-4 generated text about the one chemical, but is factual for one out of two prompts. In the case of the second compound, the model still referred to the knowledge cutoff date of 2021-09. In general, adding context did not improve neither the factuality nor specificity metrics, which was still at 70%. The quoted names of the fungus that produces the chemical were incorrect. In general, the model generated fluent, semantically coherent, highly specific text (see Fumitremorgin C in Table S.2). We observed less hallucinations compared to ChatGPT.
Llama 2: Exhibits a proclivity for generating inaccurate comparisons, particularly evident in chemical contexts. Llama 2-70B model produced 20% (4/20) responses that were both specific and factual. Smaller versions of Llama2 demonstrated inferior performance. Adding context failed to enhance either factuality or specificity metrics; instead, such augmentation resulted in a noteworthy decline, approaching zero, across all tested Llama 2 models. The Llama2 models displayed a noticeable tendency to generate comparative phrases, particularly evident when tasked with chemical comparisons, frequently resorting to statements like ’100 times more toxic than’ or ’100 times more potent than’. Specifically, the Llama2-7B produced such comparative statements in 70% of cases (14/20), while the Llama2-13B and Llama2-70B both did so in 40% of cases (8/20). These comparisons were consistently inaccurate, with a significant proportion mistakenly referencing morphine (10/30 such responses). In summary, the Llama 2 models failed to generate text characterized by fluency, semantic coherence, and high specificity.
Table 3: Performance in Task1: chemical compound definition generation. Each step from the framework (see section 2.4) evaluated separately. X/20 is the number of correct responses vs total number of prompts. The bootstrapping result reported as % of total with 95% confidence interval in brackets.
Task 1: Chemical compound definition generation
simple prompts P1-P3 context-based prompts P4-P7
Model STEP 1 STEP 2 STEP 3 STEP 1 STEP 2 STEP 3
GPT-2 2/20 (0-30) 0/20 (0-0) 0/20 (0-0) 9/20 (10-70) 0/20 (0-0) 0-20 (0-0)
GPT-3 20/20 (100-100) 6/20 (0-60) 6/20 (0-60) 20/20 (100-100) 7/20 (10-60) 7/20 (10-60)
GPT-neo 8/20 (10-70) 0/20 (0-0) 0-20 (0-0) 11/20 (20-90) 0/20 (0-0) 0/20 (0-0)
Bloom 4/20 (0-40) 0/20 (0-0) 0/20 (0-0) 6/20 (10-60) 0/20 (0-0) 0/20 (0-0)
BioGPT 14/20 (40-100) 1/20 (0-20) 0/20 (0-0) 19/20 (80-100) 7/20 (10-60) 2/20 (0-30)
BioGPT-large 16/20 (50-100) 6/20 (0-60) 5/20 (0-50) 20/20 (100-100) 7/20 (10-60) 6/20 (0-60)
Galactica* 2/10 (0-50) 0/10 (0-0) 0/10 (0-0) 9/20 (70-100) 1/20 (0-30) 0/20 (0-0)
ChatGPT 20/20 (100-100) 12/20 (30-90) 12/20 (30-90) 20/20 (100-100) 11/20 (30-80) 10/20 (20-80)
GPT-4 20/20 (100-100) 14/20 (40-100) 14/20 (40-100) 20/20 (100-100) 14/20 (40-90) 14/20 (40-100)
Llama 2-7B 16/20 (50-100) 2/20 (0-30) 2/20 (0-30) 4/20 (0-40) 0/20 (0-0) 0/20 (0-0)
Llama 2-13B 12/20 (30-90) 2/20 (0-30) 2/20 (0-30) 9/20 (20-80) 0/20 (0-0) 0/20 (0-0)
Llama 2-70B 13/20 (40-90) 4/20 (0-50) 4/20 (0-50) 12/20 (30-90) 1/20 (0-20) 1/20 (0-20)
    *Evaluation for Galactica included P3 for simple prompts and P6-P7 for context-based prompts.
4.2 Chemical-fungus relation determination via entity generation
The aim of the task was to check whether the model generates a factual relation between the chemical compound and the fungus, which is producer of the given compound. We investigated ten compounds (five with and five without antibiotic activity), the same as in Task 1 described in section 3.1. We followed the proposed evaluation framework (Figure 3): first investigated the fluency, prompt-alignment and semantic coherence of the generated output, then the relation with generated fungus name and the factuality and specificity of the whole generated text. As prompts, we provided simple clausal fragments (e.g. ‘{entity 1} is isolated from’) or with an additional noun (e.g. ‘{entity 1} is produced by fungus’) as prefix (see Table 2). All models, except for Galactica, were evaluated via the 11 prompts designs. We report aggregated results for all 110 prompts (Table 4) and an optimal prompt with the best performance (Table 5). The optimal prompt configuration varies across models. In total we manually evaluated 1440 outputs in this task.
GTP-2, GPT-neo and Bloom fail to establish chemical-fungus relation. The GPT-2 model did not generate any answer containing the name of a potential fungus from which a given chemical compound can be isolated, regardless of the designed prompt. In addition, the model hallucinated new bacteria names (e.g. Streptococcus aureus). The GPT-neo model generated only 3.6% (4/110) of correct chemical compound-fungus relations. One correct answer each for Ergosterol, Alternariol, Atromentin and Flavasperone for three different prompts. The optimal prompt produced five chemical-fungus names relations, where only two relations were correct. Out of four correct chemical compound-fungus relations, only two had factual knowledge for the entire generated text (1.8% (2/110)).
The Bloom model generated only 2.7% (3/110) of factual chemical compound-fungus relations. In general, only 11% (12/110) of answers contained a fungus name (both factual and no factual relations). The model generated answers with the name of the fungus only for the two prompts, that hint towards fungus context: ‘{entity 1} is produced by the fungus’ and ‘{entity 1} is isolated from the fungus’. Considering the optimal prompt, Bloom produced seven fungi, two of them were in factual relation with the chemical. Out of three factual chemical compound-fungus relations, only two had factual knowledge for the entire generated text (1.8% (2/110)).
GPT-3 produces at least one factual relation for 6/10 compounds. GPT-3 generated 51% (56/110) answers containing a fungus name, where 23.6% (26/110) were factual compound-fungus relations, but only in 13.6% (15/110) the whole relation description was factual. The optimal prompt resulted in five factual relations, each containing individual fungus, of which only four answers were factual for whole description. The model produced a factual relation for at least one prompt for 6/10 tested compounds. Alternariol turned out to be a chemical with the most factual relations (8 out of 10 tested prompts), which aligns with the results from section 3.1 for GPT-3 (Table S.4).
BioGPT biased towards Aspergillus, limited proportion of factual relations. BioGPT generated factual relations with fungi names for only 13.6% (15/110) of the cases, of which only ten answers had factual knowledge. Only three compounds were factually associated with the corresponding fungus from which they were isolated (Alternariol - six answers with factual relation, Atromentin - one answer with factual relation, Ergosterol - six answers with factual relation). Out of 62 generated answers that contained a fungus, 45 were Aspergillus, of which only 11 answers were factual. The two optimal prompts for this model were at 30% of the factual chemical compound-fungus relations.
BioGPT-large: at least one factual chemical-fungus relation for 9/10 compounds. BioGPT-large generated 40.0% (44/110) of the factual chemical compound-fungus relations, of which less than half (18/44) were factual. The model generated at least one factual answer for 9 out of 10 tested compounds. This implies that it is possible to achieve 90% of factual relations if picking the right prompt for each chemical, showing the dependency on prompt optimisation. The model gave the most factual answers for Ergosterol (8 out of 10 tested prompts). If an optimal prompt is considered, BioGPT-large obtained 60% factual relations. The lack of factual knowledge resulted mainly from the incorrect name of the inhibitor of a given compound. Contrary to previous models, we observed that BioGPT-large did not generate additional context, when unclear (e.g. the names of the discoverers or the year of discovery of the compound if not known) reducing the overall number of hallucinations.
Galactica biased towards a large genus Aspergillus and poor performance. Galactica generated factual chemical-fungi relations for 6 out of 10 compounds. However, for four prompts we naively accepted a genus name Aspergillus as the factual answer, despite not being a specific name (Table S.5). Only for Fumitremorgin C the answer contained the species name: Aspergillus fumigatus. Aspergillus is at the same time the most frequently generated fungus (7 times). For Aphidicolin and Myriocin the model generated no answer. The generated answers consisted only of the name of the fungus, so no factual knowledge in the entire generated text was assessed.
ChatGPT: coherent answers, regardless the prompt. ChatGPT generated 44.5% (49/110) factual chemical compound-fungus relations. The model provided all 11 fungal-name responses for eight compounds. The exception was one prompt for Conidiogenone and seven prompts for Ergosterol. The rest of the answers given for Ergosterol (4) were 100% factual with fungus relation. For three compounds all answers were factual (100% for Fumitremorgin C, Alternariol, Verrucarin A (Table S.6)) and for four compounds all prompts led to incorrect relations. The model produced very coherent answers regardless of the prompt, but only 31.8% (35/110) of them had a factuality for the whole generated description. The lack of factual knowledge was mainly due to the incorrect statement of the year of discovery/isolation of a given compound, the name of the discoverer or the class of the compound. The optimal prompt turned out to be ‘{entity 1} originally isolated from’, with 6/10 factual relations, of which the answer for only one relation had a factuality of whole answer. Almost 70% of the generated descriptions contained hallucinations.
GPT-4 gives either the factual chemical-fungus relations or no answer at all. GPT-4 generated 50.9% (56/110) factual chemical compound-fungi relations. The model produced very coherent answers regardless of the prompt design. For four compounds, all answers were factual (11/11, 100% for Fumitremorgin C, Alternariol, Verrucarin A, Myriocin). The model for all prompts for Ergosterol did not generate any relations with the fungus. The model factually recognized the compound by citing its characteristics, but without giving the names of the fungi that produce ergosterol. In most cases, the model creates a new first sentence without completing the sentence from the prompt. Rarely in the first sentence is the name of the fungus that produces the chemical compound. In the case of 13.6% (15/110) responses, the model asked for more context and/or indicated that the question is incomplete. For 4.5% (5/110) prompts the model referred to the knowledge cutoff date of 2021-09, thus not providing an answer. Three prompts were optimal with 6/10 factual relations. Out of 56 factual chemical-fungi relations, the model generated 48 answers that had factual knowledge for the entire generated text (43.6%, 48/110). The hallucination (lack of factual knowledge) was observed for  56% of outputs, mainly as an incorrect statement of the year of discovery/isolation of a given compound or the incorrect citation of the name of the synonym of the fungus.
Llama 2: the factuality increases with the model size. As the model size increases, the Llama 2 model exhibits a heightened capacity to generate fungal names and factual chemical compound-fungus relations, constituting 28.2% (31/110) of correct relations for Llama 2-70B. Notably, the Llama 2 70B accurately generated complete definitions, inclusive of the fungus name, with a success rate of 12.7% (14/110). Irrespective of the model size, Llama 2 exhibited a proclivity for response repetition and looping, with frequencies of 9% for Llama 2-7B, 4% for Llama 2-13B, and 5.5% for Llama 2-70B. Additionally, the model consistently supplied names of other chemical compounds as examples of fungi (24.5% Llama 2-7B, 34.5% Llama 2-13B, 27.3% Llama 2-70B) and occasionally substituted the expected fungal name with the year of compound discovery (13.6% Llama 2-7B, 12.7% Llama 2-13B, 25.5% Llama 2-70B). The Llama 2-7B and Llama 2-13B failed to provide fungal-name responses for the Ergosterol compound. Concerning the chemical compound Verrucarin A, all Llama 2 models erroneously postulated the name of a fungus (Verrucaria) unrelated to the actual producer of this compound (Albifimbria verrucaria or Myrothecium verrucaria), with instances recorded at 36.4% (4/11) for Llama 2-7B, 45.5% (5/11) for Llama 2-13B, and 36.4% (4/11) for Llama 2-70B. Prompt P18, structured as ’{entity 1} is produced by fungi, such as,’ yielded fungal names from all model variants for 9 out of 10 compounds. However, only Llama 2-13B and Llama 2-70B demonstrated factual relations, in 22.2% (2/9) and 55.6% (5/9), respectively.
Correctly generated relations are homogeneous and prevalent.
Considering the models with outputs that contain the fungus name, in majority they recognize one specific factual relation chemical-fungus, although the chemical can be produced by several fungi (RQ1). For instance, in all factually generated relations for Fumitremorgin C the fungus was Aspergillus fumigatus or Aspergillus (39 out of 39 correct fungi name, from 66 generated fungi in total). Other chemicals: Aphidicolin $\rightarrow$ Cephalosporium aphidicola (22/25 out of 55), Verrucarin A $\rightarrow$ Myrothecium verrucaria or Myrothecium (22/22 out of 73), (+)-Chloromonilicin $\rightarrow$ Monilinia fructicola or Monilinia (11/11 out of 53), Alternariol $\rightarrow$ Alternaria alternata or Alternaria fungus (61/67 out of 76), Flavasperone $\rightarrow$ Aspergillus (15/19 out of 39), Atromentin $\rightarrow$ Aspergillus (11/11 out ouf 54), Myriocin $\rightarrow$ Isaria sinclarii (17/21 out of 57). The least recognised chemical Conidiogenone was associated with Penicillium 9 out of 10 correct outputs (out of 55 outputs with fungi associations). Factually generated relations for Ergosterol are with two fungi: Aspergillus (11/23) and Ganoderma lucidum (7/23). Interestingly, Aspergillus is output by BioGPT and Ganoderma lucidum by BioGPT-Large. Although trained on the same corpus, we hypothesise that the higher number of parameters in BioGPT-Large (347M vs 1.5B) led to a more specific relation (RQ2).
Table 4: Performance in Task 2: chemical-fungus relation determination via entity generation. Each step from the framework (see section 2.4) evaluated separately. X/110 is the number of correct responses vs total number of prompts. The bootstrapping result reported as % of total with 95% confidence interval in brackets.
Task 2: Chemical-fungus relation determination via entity generation
simple prompts + context-based prompts P8-P19
STEP 1 STEP 2A STEP 2B STEP 3
Model
occurrence of
fungus name in
the generated text
factual occurrence
of fungus name in
the generated text
GPT-2 0/110 (0-0) 0/110 (0-0) 0/110 (0-0) 0/110 (0-0) 0/110 (0-0)
GPT-3 57/110 (20-80) 56/110 (20-80) 26/110 (0-50) 15/110 (0-40) 15/110 (0-40)
GPT-neo 29/110 (0-60) 28/110 (0-50) 4/110 (0-20) 2/110 (0-10) 0/110 (0-0)
Bloom 13/110 (0-30) 12/110 (0-30) 3/110 (0-10) 2/110 (0-10) 2/110 (0-10)
BioGPT 62/110 (30-80) 57/110 (20-80) 15/110 (0-40) 10/110 (0-30) 2/110 (0-10)
BioGPT-large 84/110 (50-100) 74/110 (40-90) 44/110 (10-70) 18/110 (0-40) 14/110 (0-40)
Galactica* 9/10 (70-100) 8/10 (50-100) 6/10 (30-90) 0/10 (0) 0/10 (0)
ChatGPT 103/110 (70-100) 101/110 (70-100) 49/110 (20-70) 35/110 (10-60) 35/110 (10-60)
GPT-4 87/110 (50-100) 77/110 (40-100) 56/110 (20-80) 48 (10-70) 48 (10-70)
Llama 2-7B 69/110 (30-90) 36/110 (10-60) 7/110 (0-30) 2/110 (0-10) 2/110 (0-10)
Llama 2-13B 70/110 (30-90) 49/110 (10-80) 14/110 (0-40) 8/110 (0-30) 4/110 (0-20)
Llama 2-70B 76/110 (40-90) 62/110 (30-80) 31/110 (0-60) 14/110 (0-30) 14/110 (0-40)
    *Evaluation for Galactica included P19 prompt only.
Table 5: Performance for optimal prompts in Task 2: chemical-fungus relation determination via entity generation. For each model an optimal prompt was selected. Evaluation according to the framework (see Figure 3). X/10 is the number of correct responses vs total number of chemicals.
Model Optimal Prompt STEP 2A STEP 2B
occurrence of fungus name
in the generated text
factual occurrence of fungus
name in the generated text
GPT-3 {entity 1} is produced by the fungus … 8/10 5/10 4/10
GPT-neo {entity 1} is produced by the fungus … 5/10 2/10 0/10
Bloom {entity 1} is produced by the fungus … 7/10 2/10 1/10
BioGPT {entity 1} is produced by fungi, such as … 10/10 3/10 1/10
{entity 1} is isolated from fungi, such as … 10/10 3/10 2/10
BioGPT-large {entity 1} is produced by the fungus … 10/10 6/10 2/10
{entity 1} is isolated from the fungus … 10/10 6/10 3/10
ChatGPT {entity 1} originally isolated from … 10/10 6/10 1/10
GPT-4 {entity 1} is isolated from fungi … 9/10 6/10 5/10
{entity 1} is isolated from fungi, such as … 9/10 6/10 5/10
Llama 2-7B {entity 1} is produced by fungi, such as … 9/10 0/10 0/10
Llama 2-13B {entity 1} is produced by fungi, such as … 9/10 2/10 1/10
Llama 2-70B {entity 1} is produced by fungi, such as … 9/10 5/10 2/10
5 Discussion
5.1 LLMs’ biases towards compounds and fungi
Bias towards Aspergillus - the most cited fungus in PubMed.
We observe that the models are biased towards certain compounds (RQ1). For instance, BioGPT generates the fungus Aspergillus as the answer in most of the prompts (79%, 45/57), which is rarely correct (only 11 times; 19%) (Table S.5). We accredit this bias to the imbalance in the training corpus, as the Aspergillus is a ubiquitous fungus in home and hospital environments. The genus Aspergillus, consists of a few hundred opportunistic species found in various climatic conditions, of which in humans, A. fumigatus is the most common and life-threatening airborne pathogen, which is particularly important among immunocompromised hosts [40]. The PubMed search outputs almost 60k references related to Aspergillus. The most cited compound from our analysis Ergosterol outputs  6000, second Aphidicolin  2500. Thus, we argue Aspergillus is over-represented in the training corpus and occurs as a statistically most probable fungus to be related to chemicals.
Rare chemical compound confused with an overrepresented biological process.
Another example, Conidiogenone was consistently not recognised by the best models (GPT-4, ChatGPT, BioGPT-large, Galactica). We argue that the models confuse Conidiogenone with conidiogenesis or conidiation (RQ1). Conidiogenesis is a widespread morphogenetic process that filamentous fungi undertake for dispersion. Conidiation is also important for pathogenicity of phytopathogens. Understanding the cellular mechanism of conidiation is a highly relevant research topic, thus it has been extensively studied. Conidiogenone is associated with nine results in a PubMed search, conidiogenesis with 346 and conidiation with 6047. The strongest and widespread stimulus for conidiation among filamentous fungi is the exposure of hyphae to the air. Often the generated texts focused on this aspect of the process, suggesting that the model encoded the signal for the representation of the process instead of the chemical compound.
Output not focused on fungi despite the prompt context.
In the case of Ergosterol, which is a compound widely known for its significance in the scientific community, the LLM specifically trained for the bio-domain (BioGPT-Large) outperforms the rest (RQ3). It generated seven factual Ergosterol-fungus relations from 11 prompts (Table S.7). For the other models, the relation to fungus was either minimally addressed or not considered at all, particularly for the models that have generally demonstrated superior efficacy (i.e., GPT-4 with 0/11 and ChatGPT with 4/11 factual responses). We observe that GPT-4 and ChatGPT focus on Ergosterol’s biosynthesis in fungi and relevance as an antifungal target, rather than on the producing fungi, despite the prompt that explicitly defines such context. Importantly, these two models do not generate text as a continuation of the prompt, but they generate new paragraphs.
5.2 Limited ability of LLMs as Knowledge Bases
In this study, we provide evidence of the limitations in the determination of factual knowledge from LLMs in the text generation tasks. Best models were able to provide only 6/10 of the factual description of chemical entities, and only 5/10 of the factual description of chemical-fungus relations. Despite small sample size in our evaluation, it clearly shows that the performance is below the threshold allowed for a systematic and reliable application (RQ1, RQ4).
5.3 High dependency and variability due to prompt design choices
A desirable property of a LLM is to a consistent answer regardless of the prompts design. For instance, the prompt ‘{entity 1} is produced by fungi, such as’ should be equivalent to ‘{entity 1} is isolated from fungi, such as’ because it refers to the same biological relation. However, we observe a variation in the output as the evidence of sensitivity for prompt design (RQ1). Similarly to Masked Language Models [41], in order to fully exploit GPTs capabilities, prompt engineering is required. In our study, we observe that providing context in the prompt improves the performance. Of note, ChatGPT and GPT-4 are the least sensitive and produce new paragraphs instead of finishing the prompt (RQ3).
5.4 Llama 2, despite large size, responses lack factuality and specificity
The analysis of Llama 2’s capabilities shows that factuality improves with increased model size. An example of this correlation is Ergosterol, not recognised in Task1 by 7B and 13B versions, but correctly defined by Llama 2 70B. Yet, even the biggest Llama2 variant, did not match the factual strength of GPT-4, ChatGPT, or BioGPT-large, delivering longer, less fluent, and semantically weaker responses. In generating chemical definitions, GPT models surpassed Llama 2 with more nuanced outputs.
5.5 Best overall performance was on large corpora, large parameters and most extensive human feedback alignment
GPT-4 outperforms ChatGPT (as well the rest tested models) in terms of factuality of the answer (RQ3). It scores 70% (7/10) and 43.6% (48/110) for chemical entity recognition and chemical-fungi relation recognition, respectively. In general, the factual knowledge of the entire generated text was higher (still significantly higher in GPT-4), but the output did not contain the relation that we asked for, so they were considered not specific. The hallucinations in GPT-4 is the lowest (54.4%). Importantly, GPT-4 produced only names of fungi which actually produce chemical compounds.
5.6 Limitations
We recognise the following limitations of our study. First, LLMs are evaluated on a specific biological dataset related to chemical compounds, fungi and antibiotic activity. The models can perform differently for other biological relation. The sample size of 10 compounds is relatively small and increasing the size would lead to more representative results. However, due to a significant effort, including domain experts, required to evaluate generated texts, we argue that our findings highlight current limitations and challenges regarding LLM’s applications in the biomedical domain.
Although we evaluated 19 prompt designs, the prompt engineering could be improved with e.g. a systematic search or optimal prompt algorithms. No fine-tuning, nor few-shot learning was performed. Evaluated LLM’s are of-the-shelf models. Moreover, this analysis concentrated on comparing and contrasting the behaviour of different models with regard to their factuality for a focused task within a scientific domain of discourse. Models have a diverse set of corpora, number of parameters and design intent, and it was not the intent of this study equalise and control for these settings. Instead, this analysis aimed to sample a highly diverse set of models, in order to elicit, which of these features are relevant for factual, knowledge-based inference within LLMs.
A total of $>$ 150 000 fungi species have been described in the literature [42]. The naming of fungal species is subject to change over time [42, 43], and moreover, the number of fungal names continue to increase over time. Despite the improvements in information sharing, the changing fungal nomenclature rules [44, 45] and providing a tool for name standardisation, the existing literature still contains plenty of synonyms, isonyms, homonyms, orthographic variants and misused names that are not in accordance with the standard nomenclature system [46]. In addition, some taxon names in the databases may be variants, synonyms or invalidly published names for various reasons. This may lead to inaccurate search results and renders the assessment of factual knowledge in generated relations challenging even for the domain expert.
6 Related Work
Large Language models achieve state-of-the-art performance in many NLP related-task, sometimes outperforming previous baselines in few-shot settings [7]. However, the factuality of the knowledge that can be extracted by prompting these models has been widely discussed. Since the first demonstration by Petroni et al. [47] on Masked Language Models (MLM), several studies have examined the mechanisms and biases underlying these predictions, questioning their applicability beyond their potential as KBs.
Cao et al. [48] showed that predictions from MLM are highly prompt-biased and that, while external context apparently helps the models, its mainly via type guidance or answer leakage. In follow-up works, many domain-specific datasets and original prompting approaches were developed for factual knowledge extraction on MLM [49, 50, 51, 52, 53, 54, 16, 55]. In the biomedical area, Sung et al. [56] extended the work of Cao et al. on a new dataset, bioLAMA, confirming their observations for biological knowledge. More recently, Kalo and Fichtel [57] extended the evaluation on a novel benchmark dataset (KAMEL) and to generative Language Models from the GPT and OPT family. Si et al. [58] evaluated the generalisability, fairness, calibration, and factuality of different prompting strategies on GPT-3. According to Mallen et al. [59], it has been observed that retaining less popular factual knowledge is hard for LLM and simply increasing their size does not significantly improve their ability to remember. Razniewski et al. [60] presented qualitative arguments that implicit LMs are not suitable as substitutes for explicit KBs, but may play a major role in the augmentation and curation of KBs.
Howard et al. [61] examined the implications of generative AI for antimicrobial advisory in infection scenario-based questions and outlined the limitations of implementing ChatGPT for interventions due to deficits in situational awareness, inference, and consistency. Li et al. [62] reported that ChatGPT performs moderately or poorly in several biomedical tests and is unreliable in actual clinical use. Furthermore, implementing ChatGPT into clinical practice is associated with numerous obstacles, such as deficits in situational awareness, reasoning, and consistency [63]. While the use of natural language processing in healthcare is not novel, the introduction of GPT-4 has elicited intense debate regarding its potential opportunities and challenges in healthcare [64].
This paper builds upon findings from several key studies to address the limitations and potentials of LLMs in scientific knowledge encoding. Zhao et al.(2021) [65], Kassner, Krojer, and Schutze (2020) [66] and Wysocki et al. (2023) [16] emphasize the inconsistency of language models and their tendency to replicate prevalent responses, a form of co-occurrence bias also noted in LLMs by Kandpal et al. (2023) [67] and Kang and Choi (2023) [68]. This tendency undermines the models’ ability to generalize, a concern supported by Razeghi et al. (2022) [69], who report an accuracy disparity tied to the frequency of terms in training data. Biderman et al. (2023) [70], Power et al. (2022) [71] and Tirumala et al. (2022) [72] explore delicate balance between memorization’s beneficial role in preventing the hallucinations, versus outputting entire sequences from their training data verbatim, as a key safety concern. Wang et al. [73] and Delmas, Wysocka, and Freitas (2023) [74] remark on LLMs’ underperformance when compared to supervised settings, positing retrieval-augmentation as a potential remedy for knowledge-dense tasks. Together, these studies highlight the need for strategies to refine LLMs’ bias and factuality, enhancing their application in biomedical domain.
7 Conclusions
This work explores the potential of LLMs to encode domain-specific factual knowledge, with an emphasis on the biomedical field. To assess this, we introduced an evaluation framework aimed at mitigating the labor-intensive nature of text generation evaluation, promoting efficiency without sacrificing accuracy. Our framework scrutinizes fluency, prompt-alignment, semantic coherence, factuality, and specificity in the generated text. Is was applied to ten off-the-shelf models, focusing on two tasks: chemical compound definition generation and chemical-fungus relation determination via entity generation. The results reveal that, although recent advancements have enhanced fluency and semantic coherence in LLMs, a reliable performance in a zero-shot setting remains elusive. The factuality of the generated content, especially in biological contexts, is still lacking, with models showing a pronounced bias towards entities that are more prevalent in the training corpus. Our findings align with previous studies across various domains [69, 73, 65, 66, 67], confirming co-occurrence biases irrespective of model size. Notably, GPT-4, while outperforming others by demonstrating a form of epistemic caution, acknowledged its knowledge limitations, thereby preventing certain hallucinations. It recognized 70% of chemical compounds and produced less than half of the factual and specific relations to fungi. This compares to the Llama2 model, which recognized 20% of compounds and 12.7% of relations to fungi. Our study addresses the increasing demand for thorough assessments of LLMs in various fields, advancing beyond the standard benchmarks that are commonly employed. This research is among the first to investigate LLM performance within the biomedical domain, specifically probing whether these models simply recall information or truly comprehend the symbolic connections between chemicals and fungi when generating definitions.
Future work. As a continuation of this study, we plan to further assess LLMs in the realm of biomedical inference, keeping pace with the rapid advancements in the field by including the latest models. We plan to extent our proposed framework addressing the need for efficient qualitative and quantitative evaluation of responses generated by LLMs. Such frameworks should go beyond the well-known benchmarks and be applicable in any specific biomedical applications, such as the one exemplified in this study (chemical-fungus relation). The promising future direction is to use prompts containing biomedical texts as contexts (abstract or full texts) combined with few-shot learning or in-context learning, as well as the integration with robust, high-recall retrieval augmentation mechanisms, e.g. [74].
Acknowledgments
The authors would like to thank the support of Joël Dumoulin, Joel Rossier and Colombine Verzat.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 965397. This work was supported by the IDIAP Research Institute and has been done in collaboration with the company Inflamalps SA and is supported by the Ark Foundation. The funding bodies played no role in the design of the study, research, writing and publication of the paper.
Supplementary Information
Supplementary Tables
Table S.1: The dataset for the experiments contains ten chemical compounds (chem name): five with (antibiotic activity = 1) and five without antibiotic activity (antibiotic activity = 0). For each relation of a selected chemical compound with a fungus (fungi name) there is a number of references (nb ref) describing the relation based on PubChem.
fungi id
fungi name
entity 2
family name pubchem id
chem name
entity 1
antibiotic
activity
0/1
nb ref
Publication year of
the latest reference
211776 Aspergillus fumigatus Aspergillaceae 403923 Fumitremorgin C 0 10 2015
119834 Alternaria alternata Pleosporaceae 5359485 Alternariol 0 9 2012
148413 Ganoderma lucidum Polyporaceae 444679 Ergosterol 0 6 2009
100745 Ganoderma australe Polyporaceae 444679 Ergosterol 0 3 2004
119872 Ganoderma applanatum Polyporaceae 444679 Ergosterol 0 3 2014
504340 Ophiocordyceps sinensis Ophiocordycipitaceae 444679 Ergosterol 0 3 2018
182069 Aspergillus nidulans Aspergillaceae 444679 Ergosterol 0 3 2009
298052 Gymnopilus spectabilis Strophariaceae 444679 Ergosterol 0 2 2005
250886 Phaeolepiota aurea Agaricaceae 444679 Ergosterol 0 2 1991
315778 Hypsizygus marmoreus Agaricaceae 444679 Ergosterol 0 2 2006
236989 Monilinia fructicola Sclerotiniaceae 121225493 (+)-Chloromonilicin 0 2 2011
247956 Penicillium aurantiogriseum Aspergillaceae 10892066 Conidiogenone 0 2 2003
257047 Cephalosporium aphidicola Cordycipitaceae 457964 Aphidicolin 1 10 2004
303853 Pleospora betae Pleosporaceae 457964 Aphidicolin 1 5 2018
337192 Pleospora bjoerlingii Neocamarosporiaceae 457964 Aphidicolin 1 5 2016
284309 Aspergillus niger Aspergillaceae 5748546 Flavasperone 1 6 2004
815927 Albifimbria verrucaria Stachybotryaceae 6326658 Verrucarin A 1 5 2012
815989 Paramyrothecium roridum Stachybotryaceae 6326658 Verrucarin A 1 3 2019
245274 Myrothecium verrucaria Stachybotryaceae 6326658 Verrucarin A 1 2 2012
283055 Dendrodochium toxicum Bionectriaceae 6326658 Verrucarin A 1 2 2004
360246 Tapinella atrotomentosa Tapinellaceae 99148 Atromentin 1 2 2007
340026 Thelephora aurantiotincta Thelephoraceae 99148 Atromentin 1 2 2005
504278 Ophiocordyceps heteropoda Ophiocordycipitaceae 6438394 Myriocin 1 2 2009
Table S.2: Examples of generated text for selected models, selected compounds as input, and selected prompts. In the ’Comment’ column there are correct answers or comments by the authors referring to semantic coherence or factuality of generated text.
Model Input Prompt Generated Text Comment
STEP 2
Factuality
BioGPT Aphidicolin {entity 1} is
Aphidicolin is a potent inhibitor of DNA polymerase
alpha and beta, but not of DNA polymerase gamma.
Aphidicolin is a potent inhibitor of
both host cell DNA polymerase alpha
and herpes simplex virus (HSV)-induced
DNA polymerase but has no effect on
DNA polymerases beta and gamma
of host cells. [ref]
0
Ergosterol {entity 1} is
Ergosterol is a natural product that has been shown
to have a wide range of biological activities.
General description which was repeated
by the model several times.
1
Aphidicolin
{entity 1} is
a compound
Aphidicolin is a compound that inhibits DNA
replication by inhibiting the enzyme DNA
polymerase alpha.
1
Aphidicolin
{entity 1} is
a substance
Aphidicolin is a substance that inhibits DNA
replication by blocking the replicative DNA
polymerase.
1
Flavasperone
{entity 1} is
a compound
Flavasperone is a compound with a wide range
of pharmacological properties.
General description which was
repeated by the model several times.
1
Verrucarin A
{entity 1} is
a compound
Verrucarin A is a compound with a unique structure
and a unique mechanism of action.
General description which was
repeated by the model several times.
0
Fumitremorgin C
{entity 1} is
a compound
Fumitremorgin C is a compound with a unique
structure and a unique mode of action.
General description which was
repeated by the model several times.
0
Alternariol
{entity 1} is
a compound
Alternariol is a compound with a wide range
of biological activities.
General description which was
repeated by the model several times.
1
Fumitremorgin C
{entity 1} is
a substance
Fumitremorgin C is a substance that is produced
by the fungus Aspergillus terreus.
Fumitremorgin C is not produced
by the fungus Aspergillus terreus.
0
Ergosterol
{entity 1} is
a substance
Ergosterol is a substance that is produced by
the fungus Ergoderma lucidum.
Model generates new name of fungus.
There is no Ganoderma lucidum
in the literature. [ref]
0
Conidiogenone
{entity 1} is
a substance
Conidiogenone is a substance isolated from
the fungus Aspergillus versicolor.
Conidiogenone is not isolated from
the fungus Aspergillus versicolor.
0
(+)-Chloromonilicin
{entity 1} is
a substance
(+)-Chloromonilicin is a substance that inhibits the
growth of the fungus Fusarium oxysporum f.
(+)-Chloromonilicin doesn’t inhibit the
growth of the fungus Fusarium
oxysporum f.
0
Verrucarin A
{entity 1} is
a substance
Verrucarin A is a substance that is produced by
the fungus Aspergillus versicolor.
Verrucarin A is not produced by the
fungus Aspergillus versicolor.
0
BioGPT-Large (+)-Chloromonilicin {entity 1} is
(+)-Chloromonilicin is a new antibiotic produced
by a strain of Streptomyces. It is active against
Gram-positive bacteria and fungi. The structure
of (+) -chloromon
Chloromonilicin is a natural
product found in Alternaria
sonchi and Monilinia fructicola .
0
ChatGPT Aphidicolin {entity 1}
Aphidicolin is a naturally occurring compound that
was originally isolated from the fungus Cephalosporium
aphidicola. It is a potent inhibitor of DNA synthesis and
is commonly used as a research tool to study DNA replication.
Aphidicolin inhibits the activity of DNA polymerases,
which are the enzymes responsible for synthesizing
new DNA strands during DNA replication. By blocking
this process, aphidicolin prevents cells from dividing and
proliferating. This makes it useful for studying the effects
of DNA replication inhibition on cellular processes such
as DNA repair, cell cycle progression, and gene expression.
Aphidicolin has also been investigated as a potential
anti-cancer agent, as many cancer cells rely on rapid
DNA replication to support their growth and survival.
However, its clinical use has been limited by its toxicity
and side effects.
Overall, aphidicolin is a valuable research tool for studying
DNA replication and has potential applications in cancer
treatment, but further research is needed to better understand
its mechanisms of action and potential therapeutic uses.
1
Aphidicolin {entity 1} is
Aphidicolin is a naturally occurring compound that is
commonly used as a research tool to study DNA replication.
It is a potent inhibitor of DNA synthesis and works by
blocking the activity of DNA polymerases, which are
the enzymes responsible for synthesizing new DNA strands
during DNA replication. By inhibiting DNA replication,
aphidicolin prevents cells from dividing and proliferating,
making it useful for studying the effects of DNA replication
inhibition on cellular processes such as DNA repair, cell cycle
progression, and gene expression. Aphidicolin has also been
investigated as a potential anti-cancer agent, but its clinical
use has been limited by its toxicity and side effects.
1
Aphidicolin
{entity 1} is
a compound
Yes, aphidicolin is a compound. It is a naturally
occurring tetracyclic diterpene, which means it is
a type of organic molecule made up of four fused
rings and derived from the isoprene unit. The chemical
structure of aphidicolin consists of 18 carbon atoms,
28 hydrogen atoms, and 2 oxygen atoms.
The chemical formula was incorrect 0
Aphidicolin
{entity 1} is
a substance
Yes, aphidicolin is a substance. It is a naturally occurring
organic compound that is found in certain fungi, including
Cephalosporium aphidicola. It can also be synthesized
chemically in the laboratory. As a substance, aphidicolin
is used as a research tool to study DNA replication and
as a potential anti-cancer agent, although its clinical use
has been limited due to its toxicity and side effects.
1
Table S.3: Examples of generated text for selected input: chemical compound Myriocin and a simple prompt: ’{entity 1}’ (for Galactica: What is the {entity 1}?) for all ten models analyzed in the paper. Factuality of generated text are provided for all generated responses.
Model Generated Text
STEP 2
Factuality
GPT-2
Myriocin  (Kaihuil) In this series, we see the relationship between three main protagonists, Kaihua, Kihua and I
0
GPT-3
Myriocin is an antifungal drug used to treat fungal infections. It is derived from the fungus Streptomyces myriocin,
and is used to treat a variety of fungal infections, including Candida
1
gpt-neo Myriocin 0
bloom
Myriocin 抗肿瘤药 抗肿瘤药 抗肿瘤药 抗肿瘤药 抗肿瘤药 抗肿瘤药
抗肿瘤药 抗肿瘤
0
BioGPT
Myriocin , a novel peptide from the venom of the scorpion Leiurus quinquestriatus hebraeus.
0
BioGPT-large
Myriocin , a specific inhibitor of serine palmitoyltransferase, was used to inhibit de novo sphingolipid
synthesis. The results showed that myriocin treatment significantly reduced the levels of
1
Galactica
Question: What is the Myriocin ? Answer: Myriocin
0
ChatGPT
Myriocin is a natural product and a type of mycotoxin that is produced by several species of fungi, including
Isaria sinclairii and Myriococcum albomyces. It was first isolated from the fungus Myriococcum albomyces in
1994. Myriocin is also known as thermozymocidin, ISP-1 or Penicillium mycotoxin, and it has been shown to have
a range of biological activities. Myriocin is a potent inhibitor of serine palmitoyltransferase, which is an enzyme
involved in the synthesis of sphingolipids. Sphingolipids are important structural components of cell membranes
and play a role in signaling pathways. By inhibiting serine palmitoyltransferase, Myriocin disrupts sphingolipid
synthesis, which can lead to apoptosis (programmed cell death) in certain cells.
Myriocin has been studied for its potential use as a therapeutic agent for a variety of conditions, including cancer,
inflammation, and autoimmune diseases. It has also been investigated for its potential as an immunosuppressive
agent in the treatment of organ transplantation and autoimmune diseases. Additionally, Myriocin has been found
to exhibit antifungal and antiparasitic activity.
1
GPT-4
Myriocin, also known as ISP-1 or thermozymocidin, is a naturally occurring compound with potent
immunosuppressive and antifungal properties. It is derived from the fungus Mycelia sterilia, specifically the Isaria
sinclairii species. Myriocin is a structural analog of sphingosine, a component of sphingolipids, which are essential
components of cell membranes in eukaryotes.
Myriocin acts by inhibiting serine palmitoyltransferase (SPT), a key enzyme in the de novo biosynthesis of
sphingolipids. This inhibition disrupts the synthesis of sphingolipids, leading to alterations in cell membrane
structure and function. Consequently, myriocin has the potential to suppress various immune responses and limit
the growth of certain fungal pathogens.
1
Llama 2-7B
is a new antibiotic that has been developed to treat bacterial infections. It is effective against many types of bacteria,
including those that are resistant to other antibiotics. Myriocin works by interfering with the bacteria‚
Äôs ability to make proteins. This prevents the bacteria from growing and multiplying, which eventually kills them.
Myriocin is available in both oral and intravenous forms.
0
Llama 2-13B
is a natural product isolated from the marine sponge Mycale hentscheli. It has been shown to have antitumor
activity in vitro and in vivo, but its mechanism of action remains unknown. In this study, we investigated
the effect of myriocin on cell cycle progression and apoptosis in human breast cancer cells.
0
Llama 2-70B
(CAS 10238-79-6) World Market Report 2019 aims at providing comprehensive data on myriocin market globally
and regionally (Europe, Asia, North America, Latin America etc.). The report includes myriocin description,
covers its application areas, manufacturing methods, patterns. It captures myriocin market trends, pays close
attention to myriocin manufacturers and names suppliers. Furthermore, myriocin prices in regional markets
can be found in the report with regards to countries and companies. In addition to the above the report determines
myriocin consumers. 3. MYRIOCIN MARKET WORLDWIDE. MANUFACTURERS AND SUPPLIERS OF
MYRIOCIN (INCLUDING CONTACT DETAILS).
0
Table S.4: Generated fungus name in Task 2: chemical-fungus relation determination via entity generation task for one chemical compound Alternariol for 11 prompts (P8-P18) (for Galactica model only P19 prompt). Fungus names that are generated correctly in bold.
Prompt GPT-2 GPT-3 GPT-neo Bloom BioGPT BioGPT-Large Galactica ChatGPT GPT-4 Llama 2-7B Llama 2-13B Llama 2-70B
Question: Which fungus
produces Alternariol ?
X X X X X X Alternaria X X X X X
Alternariol is produced by Alternaria alternata Alternaria alternata Phomopsis amygdali X Alternaria
Alternaria alternata,
A. tenuissima, A. mali
Amanita phalloides,
A. phalloides var. virosa
Alternaria alternata,
A. arborescens,
A. brassicicola,
A. carotae, A. dauci,
A. longipes, A. radicina,
A. tenuissima,
A. triticina, A. zinniae
Alternaria alternata
Alternariol is produced by the fungus Alternaria alternata Penicillium acnes Alternaria alternata Phomopsis amygdali X Alternaria
Alternaria alternata,
A. tenuissima
Alternaria alternata Alternaria alternata
Alternariol is produced by fungi
Alternaria,
Fusarium,
Penicillium
Ascomycota:
Tremellomycetes
X Alternaria
Alternaria alternata,
A. tenuissima,
A. infectoria
Alternaria Alternaria Alternaria
Alternariol is produced by fungi, such as Alternaria Aspergillus Alternaria alternata Alternaria alternata X
Alternaria alternata,
A. tenuissima,
A. infectoria
Alternaria alternata,
A. tenuissima,
A. infectoria
Penicillium verrucosum,
Aspergillus versicolor
Alternaria alternata,
A. tenuissima
Alternaria alternata
Alternariol is isolated from Alternaria Phomopsis amygdali X Alternaria Alternaria Alternaria
Alternaria alternata,
A. tenuissima
Alternaria alternata,
Aureobasidium
pullulans
Alternariol originally isolated from Alternaria alternata Aspergillus terreus Phellinus linteus X Alternaria tenuissima Alternaria alternata Alternaria alternata Alternaria alternata Alternaria
Alternariol is isolated from the fungus Alternaria alternata Alternaria alternata Phomopsis amygdali X Alternaria
Alternaria alternata,
A. tenuissima
Alternaria alternata Alternaria alternata
Alternariol is isolated from fungi Alternaria X Alternaria
Alternaria alternata,
A. tenuissima
Penicillium verrucosum,
Aspergillus niger
Alternaria alternata,
A. arborescens,
A. brassicicola,
A. carotae, A. dauci,
A. eres, A. fennica,
A. iowensis, A. longipes,
A. mali, A. mirabilis,
A. radicina,
A. tenuissima,
A. triticina
Alternaria
Alternariol is isolated from fungi, such as
Alternaria alternata,
A. tenuissima, A. solani
Alternaria alternata Phomopsis X
Alternaria alternata,
A. tenuissima,
A. infectoria
Alternaria alternata,
A. solani, A. tenuissima
Penicillium verrucosum,
Aspergillus niger
Alternaria
Alternariol is a natural product found in X Alternaria Alternaria Alternaria alternata
Alternariol is a natural chemical
compound found in
X Alternaria Alternaria Alternaria alternata Alternaria
Table S.5: Performance in the Task 2: chemical-fungus relation determination via entity generation task. Evaluation according to the framework (see Figure 3). 10 chemicals, 11 prompts (P8-P18) (for Galactica only P19 prompt), 110 outputs per model in total (for Galactica 10 outputs). Analysis of the occurrence of the most common fungus in chemical compound-fungus relation Aspergillus sin tested prompts.
Model
Occurrence of any fungus name
in the answer /
tested prompts
Occurrence of Aspergillus
in the aswer /
occurrence of any fungus name
in the the answer
Factual occurrence of
Aspergillus in
the answer / occurrence
of fungus name in
the answer
Factual occurrence of
Aspergillus in
the answer / occurrence
of Aspergillus
in the answer
GPT-2 0/110 (0) 0/0 (0) 0/0 (0) 0/0 (0)
GPT-3 56/110 (51) 21/56 (37.5) 14/56 (25) 14/21 (66.7)
GPT-neo 28/110 (25.5) 11/28 (39.3) 4/28 (14.3) 4/11 (36.4)
Bloom 12/110 (11) 3/12 (25) 2/12 (16.7) 2/3 (66.7)
BioGPT 57/110 (51.8) 45/57 (78.9) 13/57 (22.8) 13/45 (28.9)
BioGPT-large 74/110 (67.3) 20/74 (27) 13/74 (17.6) 13/20 (65)
Galactica* 8/10 (80) 7/8 (87.5) 5/8 (62.5) 5/7 (71.4)
ChatGPT 101/110 (91.8) 21/101 (20.8) 12/101 (11.9) 12/21 (57.1)
GPT-4 77/110 (70) 21/77 (27.3) 15/77 (19.5) 15/21 (71.4)
Llama 2-7B 36/110 (32.7) 6/36 (16.7) 0/36 (0) 0/6 (0)
Llama 2-13B 49/110 (44.5) 11/49 (22.5) 1/49 (2) 1/11 (9.1)
Llama 2-70B 62/110 (56.4) 17/62 (27.4) 7/62 (11.3) 7/17 (41.2)
Table S.6: Generated fungus name in Task 2: chemical-fungus relation determination via entity generation task for one chemical compound Verrucarin A for 11 prompts (P8-P18) (for Galactica model only P19 prompt). Fungus names that are generated correctly in bold.
Prompt GPT-2 GPT-3 GPT-neo Bloom BioGPT BioGPT-Large Galactica ChatGPT GPT-4 Llama 2-7B Llama 2-13B Llama 2-70B
Question: Which fungus produces Verrucarin A ? X X X X X X Aspergillus X X X X X
Verrucarin A is produced by Verrucaria nigrescens Aspergillus niger Penicillium verrucosum X Myrothecium verrucaria Myrothecium verrucaria
Penicillium verrucosum,
Aspergillus niger
Verrucaria
Verrucarin A is produced by the fungus Verrucaria nigrescens Penicillium expansum Aspergillus niger Penicillium verrucosum X Myrothecium verrucaria Myrothecium verrucaria
Verrucarin A is produced by fungi Verrucaria Neurospora crassa Penicillium X Myrothecium verrucaria Myrothecium verrucaria Verrucaria Verrucaria Verrucaria
Verrucarin A is produced by fungi, such as Verrucaria nigrescens
Penicillium roqueforti,
Penicillium verrucarinum
Aspergillus terreus Penicillium verrucosum X
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium verrucaria Penicillium verrucosum Aspergillus
Verrucarin A is isolated from Verrucaria maura Penicillium verrucosum X Myrothecium verrucaria Myrothecium verrucaria Verrucaria Verrucaria Aspergillus
Verrucarin A originally isolated from Verrucaria nigrescens Penicillium verrucosum X Myrothecium verrucaria Myrothecium verrucaria Penicillium verrucarum
Verrucarin A is isolated from the fungus Verrucaria nigrescens Penicillium expansum Aspergillus versicolor Penicillium verrucosum X Myrothecium verrucaria Myrothecium verrucaria
Verrucarin A is isolated from fungi Verrucaria Aspergillus X Myrothecium verrucaria Myrothecium verrucaria Verrucaria Verrucaria Verrucaria
Verrucarin A is isolated from fungi, such as Verrucaria maura
Conchocalyx irregularis,
Aspergillus ochraceus,
Stachybotrys chartarum
Aspergillus terreus Penicillium verrucosum X
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium verrucaria Verrucaria nigrescens
Verrucarin A is a natural product found in Verrucaria Penicillium verrucosum X
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium Verrucaria
Verrucarin A is a natural chemical compound found in Verrucaria Penicillium verrucosum X
Myrothecium verrucaria,
Stachybotrys chartarum,
Fusarium oxysporum,
Alternaria alternata
Myrothecium Verrucaria Verrucaria
Table S.7: Generated fungus name in Task 2: chemical-fungus relation determination via entity generation task for one chemical compound Ergosterol for 11 prompts (P8-P18) (for Galactica model only P19 prompt). Fungus names that are generated correctly in bold.
Prompt GPT-2 GPT-3 GPT-neo Bloom BioGPT BioGPT-Large Galactica ChatGPT GPT-4 Llama 2-7B Llama 2-13B Llama 2-70B
Question: Which fungus produces Ergosterol? X X X X X X Aspergillus X X X X X
Ergosterol is produced by
Aspergillus
terreus
Ganoderma
lucidum
X
Ergosterol is produced by the fungus
Saccharomyces
cerevisiae
Aspergillus
nidulans
Aspergillus
niger
Aspergillus
terreus
Ganoderma
lucidum
X
Ergosterol is produced by fungi X
Ergosterol is produced by fungi, such as Emericella
Aspergillus
terreus
Ganoderma
lucidum
X
Aspergillus,
Penicillium,
Fusarium
Ergosterol is isolated from
Ganoderma
lucidum
X
Ergosterol originally isolated from
Ganoderma
lucidum
X
Claviceps
purpurea
Ergosterol is isolated from the fungus
Penicillium
expansum
Aspergillus
terreus
Ganoderma
lucidum
X
Claviceps
purpurea
Ergosterol is isolated from fungi Aspergillus X
Ergosterol is isolated from fungi, such as
Aspergillus
terreus
Ganoderma
lucidum
Ganoderma
applanatum
Ganoderma
neojaponicum
X
Candida
albicans
Aspergillus
fumigatus
Saccharomyces
cerevisiae
Penicillium
chrysogenum
Cordyceps
sinensis
Ergosterol is a natural product found in X
Ergosterol is a natural chemical compound
found in
X